# Automated Static Vulnerability Detection via a Holistic Neuro-symbolic Approach

### v1

@@ 4. MoCQ
We design MoCQ, a novel static neuro-symbolic system whose workflow is outlined in Figure 2. MoCQ first analyzes open knowledge such as documentation and tool implementations to extract the query DSL, and then constructs a core DSL subset (§ 4.1). MoCQ then generates vulnerability queries (§ 4.2) through a feedback loop with a trace-driven symbolic query validator to refine (§ 4.3) and optimize (§ 4.4) the queries. Finally, MoCQ outputs the generated query, which can be applied to detect new vulnerabilities. We summarize two key techniques in MoCQ that facilitate solving the aforementioned challenges.
[…]
Query DSL Extraction and Subsetting.Query DSL Extraction and SubsettingQuery DSL Extraction and Subsetting. To help generate syntactically- and semantically-correct queries, we first obtain the formal specifications of the query DSLs used by the static analysis tools, which also define the desired output structure for LLMs. MoCQ extracts the DSL specifications—including DSL grammars, data types, compatible APIs, and their functionalities—by automatically parsing the online documentation and implementations. However, directly using all extracted DSL specifications would overwhelm an LLM because of the DSL’s complexity, and would still not enable the LLM to efficiently generate valid queries. We thus propose a language subsetting technique, which selects a core subset of DSL features from the extracted full set with the help of a few real-world query examples. This is based on our observation that query DSL often contains redundant features, for example, those that can be equally expressed by other language features. Such a DSL subset could be better understood by the LLMs to significantly reduce the query generation complexity.
[…]
Iterative Feedback Loop.Iterative Feedback LoopIterative Feedback Loop. Atop the core DSL subset, MoCQ employs an LLM to generate queries and iteratively refines them based on fine-grained feedback from our symbolic query validator. Specifically, we instrument query execution runtime, which executes and analyzes the generated queries to locate the exact locations or causes of the syntax and semantic errors. Beyond that, the validator also tracks the intermediate program states on the query execution traces and validates whether the generated query could retrieve the given vulnerability examples. During such an iterative process, MoCQ is able to gradually generate valid queries that can analyze complex real-world programs. Moreover, the validator incorporates heuristics to detect potential overfitting in queries and guides the LLM to generalize them. It also instruments the LLM to refine the query to eliminate the false positives.
@@ 4.1. DSL Specification Extraction
As mentioned in § 3.2 and detailed in Appendix A, many query execution failures arise from a lack of grammar and semantic understanding. This suggests that it is difficult for LLMs to generate valid queries without domain-specific context. The simplest way to equip pretrained LLMs with such knowledge is through few-shot learning, where a few domain-specific query examples are provided to illustrate DSL usage. However, the effectiveness of this approach depends on the coverage and diversity of the examples. In practice, for DSLs with the complexity of Joern and CodeQL, the highly structured output space cannot be effectively captured by just a handful of examples. On the other hand, fine-tuning LLMs requires a large volume of task-specific training data—in this case, valid and diverse query examples—which is often impractical to collect.
We thus propose to extract the DSL specifications in a concise way and feed the DSL specifications to the pretrained LLMs to guide query generation. Our DSL specification includes both the grammar and the semantics of the query language. The former governs the code structure while the latter encapsulates the operational meaning behind query statements like the APIs or data structures accessed, parameter processing logic, and resulting runtime behaviors. Note that the DSL specifications are tool-specific and we only need to extract once per tool.
@@ 4.1.1. DSL Extraction
The extraction process targets both syntax and semantics. To this end, MoCQ first analyzes the static analysis tool’s documentation (e.g., online use instructions) and source code (e.g., API implementations and comments). Documentation typically contains structured content (e.g., HTML tables) that describes grammar, operations, and usage patterns. We first crawl the online documentation of the tools and prompt a pretrained LLM to extract grammar rules, data types, operations, and API usage. For example, Joern’s website (N/A, 2025f) lists its query operations (termed “Steps”) along with their descriptions.
Online documentation only briefly mentions the APIs’ behaviors. We seek more precise and comprehensive definitions from their code implementations, which offer ground-truth definitions through function signatures, type annotations, inline comments, etc. We first manually identify and retrieve relevant source code files from each tool’s implementation, then leverage an LLM to automatically summarize the functionality of the functions. This process is manageable and needs to be done only once per tool, although it could be automated by linking function names from the documentation with their implementations in the source code. The prompt template used for extraction is shown in Appendix E.
@@ 4.1.2. DSL Subsetting
We design a language subsetting technique to resolve the DSL complexity issue. We observe that DSL specification contains significant redundancy not strictly necessary for constructing useful queries. A smaller common subset is often sufficient for most common queries. In essence, we limit the DSL to a more tractable subset of features, rather than exposing every possible construct, for query generation. Concretely, the subsetting process operates at the granularity of APIs, which can strike a balance between expressiveness and control—it preserves flexibility for query construction while keeping the DSL surface concise and understandable. MoCQ first associates typical parameter choices with each API. It then analyzes the full set of DSL APIs, and decides and selects these important, necessary DSL APIs for common usage. This is feasible since the LLM can refer to features’ intended behaviors or functionalities collected in § 4.1.1 to evaluate if one is necessary.
We also provide a small collection of queries to help MoCQ learn the core DSL subset. Here the LLM only needs to base on them to select the API subset from the full set, which is much simpler than generating new queries using few-shot prompting with query examples. Nevertheless, if a security analyst identifies additional needs or corner cases, they can perform a one-time augmentation to expand the DSL subset.
@@ 4.1.3. DSL Prompting
Figure 3 presents a simplified grammar for Joern using BNF. It details the fundamental grammar rules, such as the cpg_start that initializes a traversal query, and how various components connect to form complete expressions. The grammar rules explicitly define the structure and permissible compositions of the DSL, guiding models to produce syntactically valid statements. For example, by following the grammar, an LLM knows to begin queries with cpg followed by appropriate method chaining through step_chain elements, effectively avoiding common syntax errors. The grammar also clarifies how to construct complex queries using operations like filter_step and complex_step, enabling more sophisticated code analysis patterns.
@@
Figure 3. BNF grammar for Joern, extracted by MoCQ.
@@ 4.2. Query Generation
The vulnerability examples play a crucial role in our design. First, they help LLMs understand the semantics and patterns associated with each vulnerability type, especially for unseen threats. Nevertheless, MoCQ goes beyond simple pattern matching through our dedicated generalization technique described in § 4.4.2. Second, MoCQ uses these examples as test cases to validate the generated queries. Specifically, as we will show in § 4.3, MoCQ executes the query and checks whether it can retrieve the corresponding examples from the application under analysis. We believe this is a reasonable and practical assumption for deploying MoCQ in real-world scenarios because vulnerability examples are often available, for example, from the CVE database. To this end, MoCQ takes a program slice for each vulnerability example—retaining the necessary code syntax and semantics for query generation.
@@ 22       $S\leftarrow null$
24             S = Execute (B, e, S) // Capture fine-grained runtime info
@@ 4.3. Iterative Query Refinement
Given the difficulty of query generation, it is almost impossible to produce a qualified query in a single LLM attempt, even when provided with the DSL grammar and semantics. Therefore, MoCQ incorporates an iterative feedback loop (shown as the for loop on lines 4-11 of Algorithm 1) to debug and refine the query using fine-grained feedback from a trace-driven symbolic query validator. This process terminates either when a valid query is generated or when the maximum attempt threshold (MaxN) is reached.
4.3.1. Reflection with Symbolic Validation
@@
In particular, the generated query ($Q^{t}_{i}$) can be represented as a sequence with $n$ blocks denoted as $[B_{1},B_{2},\ldots,B_{n}]$. The query produces a trace when executing. When applying the $Q^{t}_{i}$ to analyze the application containing the vulnerability example $e^{t}_{i}$, MoCQ records the intermediate program states for its blocks along the trace. Specifically, when the first $j$ blocks have been executed, MoCQ collects the set of in-scope variables and their runtime values by instrumenting the query execution runtime. Formula 1 formulates the program state $S_{j}$, where $var$ represents the in-scope variable till block $B_{j}$, and $val_{j}$ is the corresponding value at current time.
(1) $S_{j}=\{var=val_{j}\mid var\in B_{\leq j}\}$
To help locate the root cause of failure, we provide the LLM fine-grained program state information with annotated trace values $[S_{1},S_{2},\ldots,S_{n}]$. Our design of feedback further includes checks for syntax errors, execution exceptions, and semantics errors.
[…]
@@
Syntax Validation.Syntax ValidationSyntax Validation. Syntax validation checks if the query complies with the extracted DSL grammar. The static analysis tool itself provides a runtime for the queries and is naturally equipped with grammar validation capability. However, it is often coarse-grained without pinpointing the exact error locations. Therefore, we instrument and hook the grammar validator to obtain the corresponding grammar rule of the errors. Alternatively, one can leverage off-the-shelf grammar parser (e.g., constructing an ANTLR (Parr, 2013)).
[…]
@@
Execution Validation.Execution ValidationExecution Validation. Similarly, MoCQ refers to the collected DSL specifications for execution exceptions. There are multiple types of semantic errors such as undefined variables, missing attributes, type mismatches, or incorrect API usage. Our symbolic validator checks these behaviors and outputs the error locations. Besides merely identifying the errors, MoCQ also searches on the collected DSL semantics to suggest (potentially correct) operations by fuzzily matching the name similarities. For example, it could suggest "<operator>.assignment" as a similar argument for "assignment" to fix the exactName exception mentioned in § 3.2.
[…]
@@
Semantic Validation.Semantic ValidationSemantic Validation. MoCQ also evaluates and inspects the semantic behaviors of the query execution. Specifically, MoCQ debugs the runtime behavior of the query with the account of the vulnerability example $e^{t}_{i}$. This leverages the knowledge of the vulnerability example to provide feedback to the LLM for refining the query. A naive approach is to check whether the query as a whole retrieves the vulnerability example (i.e., boolean feedback). However, this is insufficient for helping an LLM pinpoint the root cause when the query fails to retrieve the example. MoCQ thus filters in only the program states corresponding to the vulnerability example from the whole-project analysis, enabling the LLM to cross-check query behavior and self-correct potential issues.
@@ 4.4.1. Eliminating FPs for Precision
When using the query $Q^{t}_{i}$ to analyze the project containing the input vulnerability example $e^{t}_{i}$, it might return false positives. A false positive must be introduced at a certain code block $B_{j}$. MoCQ thus similarly leverages the program state $S_{j}$ to help the LLM reason about how the false positive is introduced. The elimination also goes through multiple iterations. Note that it is common for a static analysis query to produce false positives—many of them cannot be excluded, for example, those involving complex path constraints.
We currently assume the cases outside the collected known true positives are false positives. In our evaluation, we take a best-effort manner to manually collect all true-positive vulnerabilities in a project as the example set $E^{t}$, therefore, all cases outside are considered false positives. In practice, non-vulnerable code is much more prevalent than vulnerable code, so a report from an underdevelopment query is more like to be a false positive. Recent research has shown it is possible to leverage an LLM to assess if a case is false positive or not (Li et al., 2025). We leave this as a future work.
@@ 4.4.2. Generalization for Recall
The generated queries could be overly specific, especially when given input vulnerability examples. To mitigate the overfitting issue, MoCQ generalizes the queries to not just focus on a specific example. This is especially important to make the queries capable of real-world detection later on. As an initial step, we design a few heuristics to assess if a query is overfitting. First, MoCQ checks whether the query relies on exact constant values (e.g., hardcoded variable names and string literals) that are unique to a specific example. Such reliance limits the applicability of the query to other contexts. Second, MoCQ examines whether the query includes structural patterns or constraints that are overly specific to the layout of a particular program, such as deep AST chains or unique call sequences. Once any such situation is found, MoCQ employs the LLM to paraphrase or abstract queries by asking it to identify the core intention behind the original query and express it in a more general form to resolve the identified overfitting situation. Such generalization improves the reusability of queries and enables the detection of vulnerability variants.
@@ 4.4.3. Merging for Efficiency
After generating individual queries for different vulnerability examples, MoCQ performs query merging to consolidate detection rules and improve efficiency. A naive approach would simply apply a logical OR operation to combine all per-example queries—sequentially running each one independently. However, this can lead to performance degradation, as many queries share redundant and repetitive conditions. To address that, MoCQ introduces an LLM-assisted query merging stage. All individual queries $Q^{t}_{i}$ are given to an LLM, which attempts to merge them into a single optimized query ($Q^{t}$) while preserving their detection capability. The merged query is passed through the symbolic query validator for all vulnerability examples to verify its correctness and effectiveness (this process is omitted from Algorithm 1). If the merged query fails validation on any of the examples, MoCQ further undergoes the feedback loop to iteratively improve the merging process. By optimizing queries before real-world adaption, MoCQ minimizes redundant computation and ensures that vulnerability detection remains scalable and precise.
@@ 5. Implementation
We implemented our solution for Joern (Yamaguchi et al., 2014) and CodeQL (CodeQL, 2024) and present important implementation details in this section.

### v3

@@ 3. MoCQ
We design MoCQ, a novel static neuro-symbolic system. Its workflow is outlined in Figure 2. MoCQ first analyzes open knowledge such as documentation and tool implementations to extract the query DSL, and constructs a core DSL subset (§ 3.1). MoCQ then generates vulnerability queries (§ 3.2) through a feedback loop, with a trace-driven symbolic query validator to refine (§ 3.3) and optimize (§ 3.4) the queries. Finally, MoCQ outputs the generated query, which can be applied to detect new vulnerabilities. We summarize two key techniques that solve the aforementioned challenges.
Query DSL Extraction and Subsetting. To help generate syntactically- and semantically-correct queries, we first obtain the formal specifications of the query DSLs used by the static analysis tools, which also define the desired output structure for LLMs. MoCQ extracts the DSL specifications—including DSL grammars, data types, compatible APIs, and their functionalities—by automatically parsing the online documentation and implementations. However, directly using all extracted DSL specifications would overwhelm the LLM because of the DSL’s complexity, and would still not enable the LLM to efficiently generate valid queries. We thus propose a language subsetting technique, which selects a core subset of DSL features from the extracted full set. This is based on our observation that query DSL often contains redundant features, for example, those that can be equally expressed by other language features. Such a DSL subset could be better understood by the LLMs to significantly reduce the query generation complexity.
Iterative Feedback Loop. Atop the core DSL subset, MoCQ employs an LLM to generate queries and iteratively refines them based on fine-grained feedback from our symbolic query validator. Specifically, we instrument query execution runtime, which executes and analyzes the generated queries to locate the exact locations or causes of the syntax and semantic errors. Beyond that, the validator also tracks the intermediate program states on the query execution traces and validates whether the generated query could retrieve the given vulnerability examples. During such an iterative process, MoCQ is able to gradually generate valid queries that can analyze complex real-world programs. Moreover, the validator incorporates heuristics to detect potential overfitting in queries and guides the LLM to generalize them. It also instruments the LLM to refine the query to eliminate the false positives.
@@ 3.1. DSL Specification Extraction
To help LLMs generate syntactically and semantically valid queries, MoCQ extracts formal specifications of the query DSLs used by static analysis tools. Rather than using few-shot learning (which provides a few query examples) or fine-tuning (which requires a large amount of labeled data), we guide the model using a concise DSL specification. This specification includes both the grammar, which defines the structural rules of the query language, and the semantics, which describe the meaning and behavior of APIs, data types, and parameters. This extraction is performed once per static analysis tool.
@@ 3.1.1. DSL Extraction
The extraction process targets both syntax and semantics. To this end, MoCQ first analyzes the static analysis tool’s documentation (e.g., online use instructions) and source code (e.g., API implementations and comments). Documentation typically contains structured content (e.g., HTML tables) that describes grammar, operations, and usage patterns. We first crawl the online documentation of the tools and prompt an LLM to extract grammar rules, data types, operations, and API usage. For example, Joern’s website (N/A, 2025f) lists the query operations (termed “Steps”) along with their descriptions.
Online documentation only briefly mentions the APIs’ behaviors. We seek more precise and comprehensive definitions from their code implementations, which offer ground-truth definitions through function signatures, type annotations, inline comments, etc. Based on the results of document analysis (e.g., keywords, API name), we prompt an LLM to retrieve relevant source code files from each tool’s implementation, then to automatically summarize the functionality of the related functions.
@@ 3.1.2. DSL Subsetting
We design a language subsetting technique to resolve the DSL complexity issue. We observe that the DSL specification contains significant redundancy not strictly necessary for constructing useful queries. A small, expressive subset is often sufficient for most vulnerability detection tasks. Instead of exposing the entire DSL, we restrict query generation to this smaller, more tractable subset, which reduces complexity and improves accuracy. Concretely, the subsetting process operates at the granularity of APIs, which can strike a balance between expressiveness and control—it preserves flexibility for query construction while keeping the DSL surface concise and understandable. MoCQ first associates typical parameter choices with each API. It then analyzes the full set of DSL APIs and prompts an LLM to decide and select these important, necessary DSL APIs for common usage. This is feasible since the LLM can refer to features’ intended behaviors or functionalities to evaluate if one is necessary.
We also provide a small collection of queries to help MoCQ learn the core DSL subset. Here, the LLM only needs to base on them to select the API subset from the full set, which is much simpler than generating new queries using few-shot prompting with query examples. Nevertheless, if a security analyst identifies additional needs or corner cases, they can perform a one-time augmentation to expand the DSL subset.
@@ 3.1.3. DSL Prompting
Figure 3 presents a simplified grammar for Joern using BNF. It details the fundamental grammar rules, such as the cpg_start that initializes a traversal query, and how various components connect to form complete expressions. The grammar rules explicitly define the structure and permissible compositions of the DSL, guiding models to produce syntactically valid statements. For example, by following the grammar, an LLM knows to begin queries with cpg followed by appropriate method chaining through step_chain elements. The grammar also clarifies how to construct complex queries using operations like filter_step and complex_step, enabling more sophisticated code analysis patterns.
@@
Figure 3. Simplified BNF grammar for Joern, extracted by MoCQ.
@@ 3.2. Query Generation
The vulnerability examples play a crucial role in our design. First, they help LLMs understand the semantics and patterns associated with each vulnerability type, especially for unseen threats. Second, MoCQ uses these examples as test cases to validate the generated queries. We believe this is a reasonable and practical assumption for deploying MoCQ in real-world scenarios because vulnerability examples are often available, for example, from the CVE database and patches. To this end, MoCQ takes a program slice for each vulnerability example—retaining the necessary code syntax and semantics for query generation.
@@ 22  $S\leftarrow null$
24     S $\leftarrow$ Execute (B, e, S) // Capture fine-grained runtime info.
@@ 3.3. Iterative Query Refinement
Given the difficulty of query generation, it is almost impossible to produce a qualified query in a single LLM attempt. Therefore, MoCQ incorporates an iterative feedback loop (shown as the for loop on lines 4-11 of Algorithm 1) to debug and refine the query using fine-grained feedback from a trace-driven symbolic query validator. This process terminates either when a valid query is generated or when the maximum attempt threshold is reached.
3.3.1. Symbolic Validation
@@
In particular, the generated query ($Q^{t}_{i}$) can be represented as a sequence with $n$ blocks denoted as $[B_{1},B_{2},\ldots,B_{n}]$. The query produces a trace when executing. When applying the $Q^{t}_{i}$ to analyze the application containing the vulnerability example $e^{t}_{i}$, MoCQ records the intermediate program states for its blocks along the trace. Specifically, when the first $j$ blocks have been executed, MoCQ collects the set of in-scope variables and their runtime values by instrumenting the query execution runtime. The corresponding program state is formulated as $S_{j}=\{(var,val_{j})\mid var\in B_{\leq j}\}$, where $var$ represents the in-scope variable till block $B_{j}$, and $val_{j}$ is the corresponding value at current time. To help locate the root cause of failure, we provide the LLM fine-grained program state information with annotated trace values $PState=[S_{1},S_{2},\ldots,S_{n}]$. Our design of feedback further includes checks for syntax errors, execution exceptions, and semantics errors.
@@
Syntax Validation. Syntax validation checks if the query complies with the extracted DSL grammar. The static analysis tool itself provides a runtime for the queries and is naturally equipped with grammar validation capability. However, it is often coarse-grained without pinpointing the exact error locations. Therefore, we instrument and hook the grammar validator to obtain the corresponding grammar rule of the errors. Alternatively, one can leverage off-the-shelf grammar parser (e.g., constructing an ANTLR (Parr, 2013)).
@@
Execution Validation. Similarly, MoCQ refers to the collected DSL specifications for execution exceptions. There are multiple types of semantic errors such as undefined variables, missing attributes, type mismatches, or incorrect API usage. Our symbolic validator checks these behaviors and outputs the error locations. Besides merely identifying the errors, MoCQ also searches on the collected DSL semantics to suggest (potentially correct) operations by fuzzily matching the name similarities. For example, it could suggest "<operator>.assignment" as a similar argument for "assignment" to fix the exactName exception mentioned in § 2.2.
@@
Semantic Validation. MoCQ also evaluates and inspects the semantic behaviors of the query execution. Specifically, MoCQ debugs the runtime behavior of the query with the account of the vulnerability example $e^{t}_{i}$. This leverages the knowledge of the vulnerability example to provide feedback to the LLM for refining the query. A straightforward approach is to check whether the query as a whole retrieves the vulnerability example (i.e., boolean feedback). However, this is insufficient for helping an LLM pinpoint the root cause when the query fails to retrieve the example. MoCQ thus filters in only the program states corresponding to the vulnerability example from the whole-project analysis, enabling the LLM to cross-check query behavior and self-correct potential issues.
@@ 3.4.1. Eliminating FPs for Precision
When using the query $Q^{t}_{i}$ to analyze the project containing the input vulnerability example $e^{t}_{i}$, it might return false positives. A false positive is often introduced at a certain code block $B_{j}$. MoCQ thus similarly leverages the program state $PState$ to help the LLM reason about how the false positive is introduced and propose fixes.
In our evaluation, we take a best-effort manner to manually collect all true-positive vulnerabilities in a project as the example set $E^{t}$, therefore, all cases outside are considered false positives. In practice, non-vulnerable code is much more prevalent than vulnerable code, so a report from an underdevelopment query is more like to be a false positive. Recent research has shown it is possible to leverage an LLM to assess if a case is false positive or not (Li et al., 2025). We leave this as a future work.
@@ 3.4.2. Generalization for Recall
To mitigate the overfitting issue, MoCQ generalizes the queries to not just focus on a specific example. This is especially important to make the queries capable of real-world detection later on. As an initial step, we design a few heuristics to assess if a query is overfitting. First, MoCQ checks whether the query relies on exact constant values (e.g., hardcoded variable names and string literals) that are unique to a specific example. Such reliance limits the applicability of the query to other contexts. Second, MoCQ examines whether the query includes structural patterns or constraints that are overly specific to the layout of a particular program, such as deep AST chains or unique call sequences. Once any such situation is found, MoCQ employs the LLM to abstract queries by asking it to express the query in a more general form to resolve the identified overfitting situation. Such generalization improves the reusability of queries and enables the detection of vulnerability variants.
@@ 3.4.3. Merging for Efficiency
After generating individual queries for different vulnerability examples, MoCQ performs query merging to consolidate detection queries and improve efficiency. A naive approach would simply apply a logical OR operation to combine all per-example queries—sequentially running each one independently. However, this can lead to performance degradation, as many queries share redundant and repetitive conditions. To address that, MoCQ introduces an LLM-assisted query merging stage. All individual queries $Q^{t}_{i}$ are given to an LLM, which attempts to merge them into a single optimized query ($Q^{t}$) while preserving their detection capability. The merged query is passed through the symbolic query validator for all vulnerability examples to verify its correctness and effectiveness (this process is omitted from Algorithm 1). If the merged query fails validation on any of the examples, MoCQ further undergoes a similar feedback loop to iteratively improve the merging process.
@@ 4. Implementation
We implemented our approach in a highly modular way and integrated it with Joern (Yamaguchi et al., 2014) and CodeQL (CodeQL, 2024a). The hardware requirements for running MoCQ align with the standard recommendations for Joern (Joern, 2024) and CodeQL (CodeQL, 2024b), typically 16 GB RAM and 4 to 8 CPU cores.
We realized the runtime program state tracking by instrumenting the query execution runtime. In particular, we enhanced the query parsing process to locate syntax errors. For the other two types, we hooked the internal language interpreter. Specifically, these query languages’ interpreters typically contain a main interpretation loop that switches over the instruction types and invokes specific handlers. We thus added additional hookers for selected handlers to collect variables and their runtime values.
5. Evaluation
We extensively evaluate MoCQ to answer the following questions.
  • RQ1: New Vulnerabilities and Patterns. How effective is MoCQ in discovering new vulnerabilities in real-world applications using new vulnerability patterns?
  • RQ2: Query Effectiveness. How effective are the MoCQ-generated queries for detecting vulnerabilities? How do they compare to related approaches?
  • RQ3: Performance and Cost. What are the performance and cost of MoCQ?
  • RQ4: Ablation Study. How does each component of MoCQ contribute to its performance?
5.1. Experimental Setup
Dataset. MoCQ is applicable to a wide range of vulnerability types and programming languages. Our evaluation focuses on representative vulnerabilities across four popular programming languages (C, Java, PHP, and JavaScript), spanning systems, enterprise, and web domains. This selection covers distinct vulnerability classes from memory safety issues to injection attacks, such as use-after-free (UAF), overflow, injection, etc. We select these in part because expert-developed queries are available, enabling direct and meaningful comparisons. For each vulnerability type, we selected projects that contain documented vulnerabilities for query generation and testing. We collected from public sources (including the CVE database, GitHub issues, and patch commits) all known vulnerabilities to facilitate evaluation of both true positives and false negatives. In total, our dataset includes 306306306 vulnerabilities. The projects have an average size of  [^14] K lines of code and are thus well suited to assess the scalability of MoCQ.
Since vulnerabilities within the same project may share similarities, using them across query generation and evaluation could lead to data leakage. We thus partition the dataset at the project level into two isolated sets with manual verification to prevent data leakage: D1 includes projects containing 212212212 vulnerabilities for query generation; D2 include projects containing 949494 vulnerabilities for effectiveness evaluation; to assess generalizability, we introduce D3, consisting of 21 latest version projects not used in D1 or D2.
Configuration of MoCQ. MoCQ is a complex system with multiple configurable components, and we will systematically investigate their impact in § 5.5. Unless otherwise noted, our default setting involves generating queries from vulnerability examples drawn from ten randomly selected projects in D1 (146 out of 212 vulnerabilities), using the Claude 3.7 Sonnet model and Joern; the model is set with a temperature of 0.2 and a maximum token size of 10,00010,00010 , 000. We ran the experiments three times with an attempt threshold of 1,0001,0001 , 000 iterations on a CPU server equipped with dual 48-thread Intel Xeon processors and 187 GB RAM.
Table 1. New vulnerabilities discovered by MoCQ.
Project Vul. Type Experts’ MoCQ Status
forkcms (N/A, 2025d) PHP SQLi Reported
SyliusCmsPlugin (N/A, 2025a) PHP XSS $\times$ Reported
parse-static-imports (N/A, 2025e) JS Cmd. $\times$ Acknowledged
es2015-proxy (N/A, 2025c) JS Proto. $\times$ Acknowledged
web-worker (N/A, 2025g) JS Proto. Reported
koa-send (N/A, 2025h) JS Proto. Acknowledged
content-type (N/A, 2025b) JS Proto. $\times$ Reported
5.2. RQ1: New Vulnerabilities and Patterns
We apply the queries generated under the default setting on the latest project dataset (D3). MoCQ-generated queries effectively discover seven zero-day vulnerabilities, shown in Table 1. For comparison, we collected existing expert-developed queries from the official tool (joe, 2025) as well as from recent publications (Li and Meng, 2021; Fröberg, 2023). On the same dataset, these queries identified three vulnerabilities (which were also detected by MoCQ-generated queries) and missed the other four. The ability to outperform expert-developed queries in new vulnerability detection demonstrated the high usability of MoCQ. We responsibly reported the vulnerabilities and to date, three have been acknowledged.
Beyond new vulnerabilities, MoCQ discovered 46 new vulnerability patterns, including both vulnerability-specific patterns and three general patterns applicable across entire programming languages, as shown in Table 2. We define an atomic operation—such as a specific API call, expression type, or data flow—that is not recognized by expert-developed queries as a new pattern, and construct a minimal PoC vulnerability example to validate each new pattern. For example, four new PHP type juggling mechanisms were found by MoCQ, including built-in functions array_search and array_flip, where implicit type casting could occur; a new general way to propagate JavaScript data flows via Reflect.set, which could benefit detection of different vulnerability types.
Although expert-developed queries for these fundamental vulnerability types had undergone multiple rounds of refinement (see § 5.4), MoCQ still uncovered critical patterns that had been overlooked. Each pattern represents a blind spot in the detection logic, and missing even one can result in an entire class of vulnerabilities going undetected. This highlights the challenge of manually designing comprehensive query coverage, even for well-known bug types. In contrast, MoCQ benefits from LLMs’ ability to capture and enumerate subtle vulnerability behaviors. We submitted the new patterns to the query maintainers, and to date, five have been merged into upstream tools’ repositories.
Table 2. Overview of 46 new vulnerability patterns uncovered by MoCQ, including 34 distinct operations in the row of phar*.
Vul. Type Pattern Functionality Status
General PHP htmlspecialchars Ctx-sensitive sanitization Reported
PHP type. in_array Implicit type casting Merged
PHP type. array_search Loose comparison Merged
PHP type. array_flip Implicit type casting Reported
PHP type. case in switch Type coercion in case matching Reported
PHP deser. *phar Object deserialization Reported
PHP deser. copy Object deserialization Reported
General JS Object.assign Data propagation Reported
General JS Reflect.set Data manipulation Merged
JS proto. Object.assign Data propagation Merged
JS proto. Object.defineProperty New property manipulation Reported
JS proto. obj.__defineGetter__ New property manipulation Merged
C/C++ UAF Reallocation via malloc Data propagation Reported
Table 3. Vulnerability detection using Joern on the testing dataset.
MoCQ Experts’ MoCQ+Experts’ Pure LLM Generation Time Analysis Time
Vul. Type TPtotal TP FP Rec. Prec. TP FP Rec. Prec. TP FP Rec. Prec. TP FP Rec. Prec. Exp.’ MoCQ Exp.’ MoCQ
PHP SQLi 19 16 14 0.84 0.53 17 21 0.89 0.45 18 23 0.95 0.44 9 45 0.47 0.17 2d 15.3h 23m 33m